OpenStudy (saranya):

You are given functions f and g such that f(n)=O(g(n)). Is f(n)∗log2(f(n)c)=O(g(n)∗log2(g(n))) ? (Here c is some constant >0. You can assume that f and g are always bigger than 1. a) true b)false c)depends on c d) depends on f and g

OpenStudy (anonymous):

I think could go this way: f(n)=O(g(n)) means |f(n)|<= M |g(n)| or |f(n)|/|g(n)| <= M so you have to prove, given that f(n)=O(g(n)), that |f(n)∗log2(f(n)c)|/|g(n)∗log2(g(n))| <= N |f(n)∗log2(f(n)c)|/|g(n)∗log2(g(n))| <= M|log2f(n)c)|/|log2(g(n))| = =M log2|(f(n)c -g(n))| <= M log2 (|M|g(n)|c - g(n)|). So i think it depends on c.

OpenStudy (anonymous):

@saranya

OpenStudy (anonymous):

i am going to guess no \[f=O(g)\iff \lim_{n\to \infty}\frac{f}{g}=M\] for some M if i remember correctly so you want \(\lim_{n\to \infty}\frac{f(n)\log(cf(n))}{g(n)\log(g(n))}\) to be finite i guess it comes down to computing \(\lim_{n\to\infty}\frac{\log(cf(n))}{\log(g(n))}\)

OpenStudy (saranya):

But acc. to Big oh Definiton it is true ryt ?

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{\log(c)+\log(f(n))}{\log(g(n))}\] so i guess as long as \(\lim_{n\to \infty} g(n)\neq 1\) you are ok hmmm

OpenStudy (anonymous):

it is early so i would make no bets on this, but i am going to say it does not depend on c of course you will want a proof, but maybe before you try one do it by example. i.e. check with some functions you know and see what happens

OpenStudy (saranya):

yea okay thanks :)

OpenStudy (anonymous):

sorry i cannot provide an actual solution good luck

OpenStudy (saranya):

no probs :)

OpenStudy (saranya):

Can anyone help out this question....? Arrange the following functions in increasing order of growth rate. ie. if g(n) follows f(n) in your list, then f(n) is necessarily O(g(n)). a)2^2n b)2^n^2 c)n^2log(n) d)n e)n^2n

OpenStudy (saranya):

@ satellite can yu reply m for this ??

OpenStudy (da_scienceman):

f cannot be same as g since from the defn of f(n) = O(g), f != g

OpenStudy (da_scienceman):

but one can as well argue that since both f and g have no relationship but all positive, it is pointless to compare both in that sense....u should revise definitions of big O and their relation in order to argue properly